OpenStudy (anonymous):

Find all solutions by factoring: t^2+9t+18=0

OpenStudy (anonymous):

What two numbers when added to together equal 9 and when multiplies together equal 18?

OpenStudy (anonymous):

3 and 6

OpenStudy (anonymous):

So to get t^2 what do you have to do?

OpenStudy (anonymous):

(t+3)(t+6)=0

OpenStudy (anonymous):

Very good..now what are the values for t going to be?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

(t+3)(t+6)=0. When you factor an equation that equals 0 you then take each factor and solve for 0 to get your possible solutions..ie t+3=0..What would the value for (t) be?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so are the solutions: (0,3,6)

OpenStudy (anonymous):

No. You have factored the equation correctly when you came up with (t+3)(t+6) =0. So the value for t in the equation t+3=0 we solve by subtracting 3 from both sides and we get t+3-3=0-3 or t=-3. So one solution for t is (-3). Now what would be the other solution?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

No. Look at it this way in the equation (t+3)(t+6)=0 We are multiplying correct. So if we multiply anything by 0 we get 0. So by making t = -3 the equation becomes (-3+3) (-3+6) = 0 or 0 times 3 which equals 0. Do you see this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In factoring when the results equal zero all we need to do is take each factor and solve it for 0. So t+3 = 0 then t has to be -3 So t+6 =0 then t has to be ?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

so the solution set is (-3, -6)

OpenStudy (anonymous):

You got it

OpenStudy (anonymous):

thank you